there are two tangents drwn from point P to parabola Y^2=4AX if m2=2m1 find the locus of point P

Grade:12th pass

1 Answers

Devki Nandan Malav
6 years ago
Remember the equation of tangent in slope form of standard parabola y2=4ax i.e y=mx+a/m
Let the locus point is (h,k). Beacuse this point lie on the tangent  so it satisfy the equation and we get 
K=mh+a/m 
m2h-km+a=0 also given that m2=2m1
Read sum of roots i.e. m1+m2=k/m so we get   m1=k/3m and this is a root of the quadratic so put in quadratic we will get the required locus